How many solutions does 5(x -4) = 5(x-9) have?
Answer by KMST(5328)

None.
5(x-4)=5(x-9) happens when, and only when
x-4=x-9.
Those 2 equation are equivalent, meaning that they have exactly the same solutions.
Then, x-4=x-9 happens wben, and only when
-4 = -9 .
That happens never.
There is no value of x that will make
4 = 9 , or -4 = -9 .